A tugboat goes 120 miles upstream in 20 hrs. the return trip downstream takes 5 hrs. find the speed of the tugboat without a current and the speed of the current
The speed of the tugboat is
The speed of the current is
simplify your answers

Answer by ankor@dixie-net.com(22740)

): A tugboat goes 120 miles upstream in 20 hrs.
the return trip downstream takes 5 hrs.
find the speed of the tugboat without a current and the speed of the current
:
Let s = boat speed in still water
Let c = speed of the current
then
(s-c) = effective speed upstream
and
(s+c) = effective speed downstream
:
Write distance equation for each way. (dist = time * speed
20(s-c) = 120
5(s+c) = 120
Simplify both equations, divide the 1st equation by 20, divide the 2nd equation by 5, results:
s - c = 6
s + c = 24
-----------Adding eliminates c find s
2s = 30
s = 30/2
s = 15 mph boat speed in still water
:
You can find the current, check the solutions in the original equations
: